Combined Probabilities (Cambridge International IGCSE Mathematics)
Combined Probability
How do I calculate combined probabilities?
-
You can calculate probabilities of one event after another without needing tree diagrams
- These are called combined (or successive) probabilities
-
There are two rules to learn
- And means multiply and or means add
- P(A and B) = P(A) x P(B)
- P(AA or BB) = P(AA) + P(BB)
-
Try to rephrase each question using and / or
-
For example, when flipping a coin twice:
- P(two heads) = P(head and head)
- P(both the same) = P(head and head or tail and tail) = P(HH) + P(TT)
-
-
Remember that P(not A) = 1 - P(A)
What does independent mean?
-
Independent events are events that do not affect each other
- e.g. the probability of rolling a 6 on a fair dice and the probability of getting a head when flipping a coin
-
Be careful: questions 'without replacement' are not independent
-
e.g. the probability of taking a red card out of a pack, not replacing it, then finding the probability of taking a second red card out of the same pack
- The first event affected the number of cards left for the second event
-
Worked Example A box contains 3 blue counters and 8 red counters. A counter is taken at random and its colour is noted. The counter is put back into the box. A second counter is then taken at random, and its colour is noted.
Work out the probability that
(a) both counters are red,
Answer:
P(both red) = P(red and red) This is P(red) × P(red) using the 'and rule'
(b) the two counters are of different colours.
Answer:
P(different colours) = P(blue and red or red and blue) This is P(B and R) + P(R and B) using the 'or rule' This is P(B) × P(R) + P(R) × P(B) using the 'and rule' twice
Was this revision note helpful?
Build on this topic